OpenStudy (anonymous):

The acute angles of a right triangle are in a 3:5 ration. Find the measure of each angle.

OpenStudy (anonymous):

*ratio

OpenStudy (anonymous):

So far, I have 3x+5x+90=180

OpenStudy (anonymous):

90/8x=11.25 then plug that in to 3(11.25) and 5(11.25) and get the answer

OpenStudy (anonymous):

But is it right? I mean, last time you assumed it was a right triangle. I just want to know if you are credible >_<

OpenStudy (anonymous):

i guessed and didn't read the problem. i thought it said sides. bc a 3,4,5 triangle is a right triangle with sides. just ask yourself "does the triangle add up to 180? and is this a 5:3 ratio with the answers you get

Directrix (directrix):

So far, I have 3x+5x+90=180 8x + 90 = 180 8x = 90 x = 90/8 = 45/4 3x = 3(45/4) = 135/4 = 33 and 3/4 or 33.75 degrees 5x = 5 ( 45/4) = 225/4 = 56 and 1/4 or 56.25 degrees